The measures of the angles in a triangle are x, 2x, and 3x. The triangle is
acute
right
obtuse

Answer by Edwin McCravy(20081)

The sum of the interior angles of any triangle is 180°, therefore

 x + 2x + 3x = 180°

          6x = 180°
 
           x = 30°

Therefore the three angles are

 x = 30°
2x = 2(30°) = 60°
3x = 3(30°) = 90°

It is a right triangle because it contains a 90° angle.

Edwin